Aijia invests $2641 in a simple interest account. The annual interest rate is 3.6%.


How much interest will she have earned at the end of 5 years?


1
Expert's answer
2021-04-20T16:24:56-0400

"I=\frac{PRT}{100}=\frac{2641\times 3.6\times 5}{100}=\$475.38"